OpenStudy (anonymous):

20% of the ointment is too weak and 60% is way too strong, but 25% is just right. To get 25% you have to mix 24 fluid onces in volume. How many fluid onces of each did we use? I give out medals :)

OpenStudy (dumbcow):

A +B = 24 .2A + .6B = .25(24)

OpenStudy (dumbcow):

solve using elimination multiply bottom equation by -5 A + B = 24 -A -3B = -30 ------------- -2B = -6 B = 3 A + 3 = 24 A = 21

OpenStudy (anonymous):

can u please show me ??? I will give u a medal :)

OpenStudy (dumbcow):

?? i already wrote the solution

OpenStudy (dumbcow):

which part do you not understand